Hi.I need help on solving equations by adding or subtracting.for ex. :n+1/3=11/12 thanks.(:
Answer by mananth(16946)

n+1/3=11/12
.
-1/3 -1/3

n=11/12 -1/3
n=(11-4)/12
n=7/12
n=7/12
..
m.ananth@hotmail.ca